# Kandinsky 3.0 Technical Report

### v2

@@ 7 Limitations
Even though the current system can generate high-quality realistic images and successfully cope with diverse and complex language concepts, there are ways for further improvement. Among them is the improvement of semantic coherence between the input text and the generated image due to more efficient use of the Flan-UL2 text encoder potential. In addition, it is possible to improve the image quality based on human evaluations significantly.
@@ 9 Conclusion
In this report we highlighted the most significant advantages of our new text-to-image generative model – Kandinsky 3.0. First we described the architecture itself and its valuable blocks that lead to high-fidelity generations and better quality metrics values. Further we show several main applications of Kandinsky 3.0 and the quality of inpainting/outpainting generation shows nearly perfect result in comparison with previous experiments in terms of naturalness of a missing area filling. Improving the text encoder size and extending the main diffusion U-Net we achieved higher human evaluation scores in comparison with Kandinsky 2.2. It should be mentioned that both measured quality indicators – text understanding and visual quality improved. Comparing with DALL-E 3 we are not so good for both indicators, whereas for SDXL we achieved significantly higher scores in both experiments. We also used Kandinsky 3.0 as a backbone for our text-to-video end-to-end generation model. In terms of our future research challenges still remain high-fidelity text generation and photorealistic faces, physics-controlled scene (lighting, positioning, focus and exposure, etc.) and text encoder research for better details understanding.
Appendix A. Acknowledgements
@@
  • Sber AI Research team: Sofia Kirillova, Arseniy Shakhmatov, Mikhail Shoytov, Sergey Nesteruk, Ilya Ryabov, Mikhail Martynov, Anastasia Lysenko, Zein Shaheen;
  • Anton Razzhigaev and Elizaveta Dakhova from AIRI;
@@
  • Tatyana Paskova, Angelina Kuts and prompt engineering team.
@@
Appendix B. Additional Generation Examples
@@
Figure 7: Example images from side-by-side comparison. Prompt: A beautiful girl.
Figure 8: Example images from side-by-side comparison. Prompt: Barbie and Ken are shopping.
Figure 9: Example images from side-by-side comparison. Prompt: A bear in Russian national hat with a balalaika.
Human evaluation results

### v3

@@ 8 Limitations
Even though the current system can generate high-quality realistic images and successfully cope with diverse and complex language concepts, there are ways for further improvement. Among them is the improvement of semantic coherence between the input text and the generated image due to more efficient use of the text encoder potential. Challenges still remain high-fidelity text generation and photorealistic faces, and physics-controlled scene (lighting, positioning, focus and exposure, etc.).
@@ 10 Conclusion
In this report we highlighted the most significant advantages of our new text-to-image generative model – Kandinsky 3.0. Improving the text encoder size and extending the main diffusion U-Net we achieved higher human evaluation scores in comparison with Kandinsky 2.2. It should be mentioned that both measured quality indicators – text understanding and visual quality improved. Comparing with SDXL we achieved significantly higher scores for both indicators.
We described the acceleration of our model using distillation (Kandinsky 3.1 model). This allowed us to increase the inference speed by 20 times and reduce the number of steps in the reverse diffusion process to 4 without decrease in quality. We also show that the Kandinsky 3.0 model can be successfully used in various applications – inpainting/outpainting, image editing, image-to-video, and text-to-video.
A. Acknowledgements
@@
  • Sber AI Research team: Nikolay Gerasimenko, Elizaveta Dakhova, Sofia Kirillova, Mikhail Shoytov, Zein Shaheen, Anastasia Yaschenko;
  • Anton Razzhigaev from FusionBrain research team at AIRI;
@@
  • Tatyana Nikulina, Angelina Kuts, Anton Bukashkin and prompt engineering team;
  • Arseniy Shakhmatov, Anastasia Lysenko, Sergey Nesteruk, Ilya Ryabov and Mikhail Martynov (ex-Sber AI).
@@
B. Generation Examples
@@
Distillation
Figure 8: Prompt: Cute red lion cub in a bath with foam.
Figure 9: Prompt: Close-up of a little boy in a tuxedo with a bouquet of white daisies, watercolor painting.
Figure 10: Prompt: Huge hot air balloon soars in the fabulous steppes and mountain ranges, freshness, high image quality.
Figure 11: Prompt: Cute little panda stands in a kimono, a very kind cute look, cute eyes, kindness.
Prompt beautification
Figure 12: Prompt: Close-up photo of a beautiful oriental woman, elegant hijab-adorned with hints of modern vintage style. Without/With LLM.
Figure 13: Prompt: A hut on chicken legs. Without/With LLM.
Figure 14: Prompt: Lego figure at the waterfall. Without/With LLM.
@@
Image Editing with IP-Adapter
Figure 16:
Figure 17:
Figure 18:
Figure 19:
@@
Figure 21: Prompt: Tomatoes on a table, against the backdrop of nature, a still life painting depicted in a hyper realistic style.
Figure 22: Prompt: Photo of Russian apartment buildings with many windows at night, lights on inside each window, the building is lighted from behind by warm yellow and white lights, it’s a wide shot showing an entire multistorey block of flats, there are more houses in front of us, high resolution, hyper realistic.
Figure 23: Prompt: Funny cute wet kitten sitting in a basin with soap foam, soap bubbles around, photography.
Figure 24: Prompt: Maximally realistic portrait of a jolly old gray-haired Negro Rastaman with wrinkles around his eyes and a crooked nose in motley clothes.
C. Side-by-side human evaluation results
Human evaluation results for text-to-image generation
@@ C. Side-by-side human evaluation results
Human evaluation results for distillation
Figure 34: Overall results of side-by-side human comparison between Kandinsky 3.0 and Kandinsky 3.0 Distilled (Kandinsky 3.1).
Figure 35: Results of side-by-side human comparison between Kandinsky 3.0 and
Kandinsky 3.0 Distilled (Kandinsky 3.1) for text comprehension.
Figure 36: Results of side-by-side human comparison between Kandinsky 3.0 and
Kandinsky 3.0 Distilled (Kandinsky 3.1) for visual quality.
Human evaluation results for prompt beautification
Figure 37: Overall results of side-by-side human comparison for Kandinsky 3.0 with and without prompt beautification.
Figure 38: Overall results of side-by-side human comparison for Kandinsky 3.0 with and without prompt beautification for text comprehension.
Figure 39: Overall results of side-by-side human comparison for Kandinsky 3.0 with and without prompt beautification for visual quality.
Figure 40: Overall results of side-by-side human comparison for Kandinsky 3.0 Distilled (Kandinsky 3.1) with and without prompt beautification.
Figure 41: Overall results of side-by-side human comparison for Kandinsky 3.0 Distilled (Kandinsky 3.1) with and without prompt beautification for text comprehension.
Figure 42: Overall results of side-by-side human comparison for Kandinsky 3.0 Distilled (Kandinsky 3.1) with and without prompt beautification for visual quality.